Hi i'm having a problem with this question
3(x+2)-2(3x-5)=10
hope you can help me out

Answer by venugopalramana(3286)

3(x+2)-2(3x-5)=10
3x+6-6x+10=10
-3x=10-10-6=-6
x=-6/-3=2